Lauras Bakery Beginning Work-in-Process Inventory Direct Materials Used Direct Labor Manufacturing Overhead Total Manufacturing Costs Incurred 14,500 10,300 45,400 during the Year Total Manufacturing Costs to Account For Ending Work-in-Process Inventory Cost of Goods Manufactured 56,000 51,600
Answer #1

Solution

Working

Laura's bakery

(A)

Beginning work in progress Inventory

$ 10,600.00

(B)

Direct material used

$ 14,500.00

(C )

Direct labor

$ 10,300.00

(D)

Manufacturing Overheads

$ 20,600.00

(E)=(B+C+D)

Total Manufacturing cost Incurred during the year

$ 45,400.00

(F)=(E+A)

Total Manufacturing cost to Account For

$ 56,000.00

(G)

Ending work in progress Inventory

$    4,400.00

(H)=(F-G)

Cost of Goods manufactured

$ 51,600.00

Working Note 1

Total Manufacturing cost Incurred during the year=

(Direct Material+ Direct Labor+ Manufacturing Overheads)

If we change the equation we get    Manufacturing Overheads=

(Total Manufacturing Cost during the year-Direct Material -Direct Labor)

Manufacturing Overheads=

(45400-14500-10300)

Manufacturing Overheads=

$ 20,600.00

Working note 2

Total Manufacturing cost to Account For=

(Total manufacturing cost incurred during the year +Beginning WIP )

If we change the equation we get     Beginning WIP=

(Total Manufacturing cost to Account For-Total Manufacturing Cost Incurred During the year)

Beginning WIP=

(56000-45400)

Beginning WIP=

$ 10,600.00

Ending WIP =

(Total Manufacturing cost to Account For-Cost of Goods manufactured)

Ending WIP =

(56000-51600)

Ending WIP =

$    4,400.00